# Reinforcement Twinning: from digital twins to model-based reinforcement learning

### v3

@@ 5.1 Wind Turbine Control
The wind energy community has put significant effort into the development of digital twins of onshore and offshore wind turbines, with applications ranging from load estimations (Pimenta et al., 2020; Branlard et al., 2023) to real-time condition monitoring (Olatunji et al., 2021; Fahim et al., 2022). In this work, we solely focus on the rotor’s aerodynamic performance and the twinning process seeks to identify a closure law for the power coefficient from data, starting from (25) and (26). In fact, the BEMT, commonly used to obtain the $C_{p}$ curve, assumes that the flow is steady, and this is questionable in the presence of large fluctuations in the wind speed (Leishman, 2002). Moreover, the coefficient can change over time as a result of turbine wear (e.g. due to blade erosion) or in the presence of extreme conditions (e.g. ice, dirt or bug buildup) (Staffell and Green, 2014). This is why any strategy based on a pre-computed $C_{p}$ curve is likely sub-optimal (Johnson, 2004, 2006) and the idea of continuous updates of this closure law from operational data has merit. Although this exercise is purely illustrative, the approach can be readily extended to more complex and multi-disciplinary model formulations.
The digital twin employs the power coefficient parametrization formulated by Saint-Drenan et al. (2020)
@@
$\mathcal{J}_{p}=\frac{1}{2}\int_{0}^{T}\Big{(}\omega(t;\bm{w}_{p})-\check{\omega}(t)\Big{)}^{2}\;\mathrm{dt}\,.$ (28)
@@ 5.2 Trajectory control of flapping wings micro air vehicles
This test case considers the control of a Flapping-Wing Micro Air Vehicle (FWMAV) which aims to move from one position to another in the shortest possible time. FWMAVs are sometimes considered better alternatives to fixed-wing and rotary-wing micro aerial vehicles because of their higher agility and manoeuvrability (Haider et al., 2020). The design and control of these drones seek to mimic the remarkable flight performances of hummingbirds, which can undertake complex manoeuvres with an impressive response time (see, for example, Cheng et al. 2016 and Ortega-Jiménez and Dudley 2018).
Reproducing these performances requires advanced knowledge of the flapping wing’s aerodynamics in the low Reynolds number regime, as well as the fluid-structure interaction resulting from the wing’s flexibility. These remain significant challenges for advanced numerical simulations (Fei et al., 2019; Xue et al., 2023). The proposed test case shows how simplified models of the wing’s aerodynamics could be derived in real-time while an FWAVs seeks to achieve its control target. The considered problem is illustrated in Figure 5(a). A drone initially at $\bm{x}_{0}=[x_{0},z_{0}]$ m and $\bm{\dot{x}}_{0}=0$ m/s is requested to move to $\bm{\tilde{x}_{f}}=[\tilde{x}_{f},\tilde{z}_{f}]$ m as fast as possible, and then hover in that position (i.e. $\dot{\bm{\tilde{x}}}_{f}=[0,0]$) by continuously adapting its flapping wing motion. We challenge the controller by introducing gusty wind disturbances $\bm{u}_{\infty}$,
@@
The body frame is first pitched by a stroke plane angle $\beta_{w}(t)$ along $y^{\prime}$ to define the stroke plane frame ($x^{\prime\prime},y^{\prime\prime},z^{\prime\prime}$). The wing tips are constrained to flap in the stroke plane ($x^{\prime\prime},y^{\prime\prime}$) as it flaps along the normal $z^{\prime\prime}$. The flapping frame ($x^{\prime\prime\prime},y^{\prime\prime\prime},z^{\prime\prime\prime}$) follows this rotational motion and defines the flapping angle $\phi(t)$ between the wing symmetry axis $y^{\prime\prime\prime}$ and the lateral direction $y^{\prime\prime}$ (Figure 5(b)). The last frame is the wing-fixed frame $(\xi,\gamma,\zeta)$ that results from the pitching rotation along the wing symmetry axis. The pitching angle $\alpha(t)$ is then defined between the chord-normal direction $\xi$ and the stroke plane axis $x^{\prime\prime\prime}$. The reader is referred to Whitney and Wood (2010) and Cai et al. (2021) for a complete treatment of the flapping wing kinematics. We select an harmonic parametrization for $\phi$ and $\alpha$:
@@
We consider a drone with mass $m_{b}=3$ g and a constant flapping frequency of $f=20$ Hz. Given the initial position $\bm{x}_{o}=[0,0]$ m, we set the target position at $\bm{x}_{f}=[5,5]$ m. We assume that the manoeuvre should take less than $n_{t}=100$ flapping cycles, hence we consider an observation time of $T_{o}=5$ s. This defines the observation time for the twinning. We fix the pitching amplitude to $A_{\alpha}=45$ deg and let the flapping amplitude vary in the range $A_{\phi}\in[50;88]$ deg. These values are comparable to what is observed in hummingbird’s flight, although the harmonic parameterization in (29) oversimplifies their actual wing dynamics (Kruyt et al., 2014). The stroke plane angle is bounded in the range $\beta_{w}\in[-30,30]$ deg.
We assume that the drone’s wings are semi-elliptical and rigid with a span $R=0.05$ m and a mean chord $\overline{c}=0.01$ m. The wing roots are offset of $R_{0}=0.0225$ m from the body barycenter, which is also the centre of rotation of the wings. We decouple their motion from the body such that the wing inertia does not influence the body motion (Taha et al., 2012). This hypothesis simplifies the body dynamics, which becomes only a function of the aerodynamic forces produced by the wings and the gravitational force. These are assumed to apply on the barycenter of the body, which is treated as a material point. Therefore, restricting the dynamics to the $(x,z)$ plane, the force balance gives
@@
[…]
To compute the wing forces in (34) and (35) we follow the semi-empirical quasi-steady approach of Lee et al. (2016), based on the work by Dickinson et al. (1999). This is a common approach in control-oriented investigations (see e.g Cai et al. (2021); Fei et al. (2019)) because it provides a reasonable compromise between accuracy and computational cost, at least for the case of smooth flapping kinematics and near hovering conditions (see Lee et al. (2016)).
The approach is based on BEMT, which partitions the wing into infinitesimal elements (see Figure 6(a)). Each element produces an infinitesimal aerodynamic force that must be integrated along the span to get the total lift and drag. In the case of the smooth flapping motion (29), the forces due to the flapping motion dominate all the other physical phenomena (Lee et al. (2016)) so that the lift $L_{w}$ and drag $D_{w}$ expressions are
@@
The force coefficients $C_{L,w}$ and $C_{D,w}$ in equations (37) and (38) are modeled similarly to Lee et al. (2016) and Sane and Dickinson (2001):
@@
where $\bm{w}_{p}=[a,b,c]$ are the closure parameters for this problem. This parametrization approximates the influence of the Leading Edge Vortex (LEV), generated at a high pitching angle due to flow separation and stably attached on the suction side (see Sane (2003)). The corresponding force coefficients depend on the effective angle of attack $\alpha_{e}=\cos^{-1}(U_{w,z}$/$||\bm{U_{w}}||_{2})$, defined between the chord direction $\zeta$ and the relative velocity of the wing $\bm{U_{w}}$.
@@
In this test case, the real environment is simulated using $\bm{w_{p}}=[a,b,c]=[1.71,0.043,1.595]$ in (37) and (38), as approximated from (Lee et al., 2016) for semi-elliptical wings. For the digital twin, these are the parameters to be identified while interacting with the environment to minimize the cost function:
@@ 5.3 Thermal Management of Cryogenic Storage
We consider the thermal management of a cryogenic tank. This is a fundamental challenge in long-duration space missions, which require storing large amounts of cryogenic propellant for long periods. These fluids must be stored in the liquid phase to maximize the volumetric energy density. However, their storage temperatures are extremely low (e.g., -250 ${}^{\circ}$ for liquid hydrogen), and their latent heat of evaporation is much smaller than that of non-cryogenic propellants. These combined factors challenge the storage because evaporation due to thermal loads produces a continuous pressure rise over time (Salzman, 1996; Motil et al., 2007; Chai and Wilhite, 2014).
@@
The efficient management of this problem requires a combination of advanced insulation strategies (see Mer et al. 2016b; Jiang et al. 2021) and venting control strategies (Lin et al., 1991). In this test case, we consider a classic active Thermodynamic Venting System (TVS), similar to the one presented in Lin et al. (1991) and recently characterized in Imai et al. (2020); Qin et al. (2021). Figure 7(a) shows the schematics of the TVS configuration.
@@
Figure 8: Schematic of the thermal management problem considered in this work. A cryogenic tank of volume $V_{t}$, containing a volume $V_{l}$ of saturated liquid and $V_{t}-V_{l}$ saturated vapour, is subject to a heat load $Qh(t)$. The figure on the right traces the state of the system on the temperature-specific entropy diagram considering the simplified model used to simulate the environment.
@@
We consider a tank of volume $V_{t}$ filled with a volume of liquid $V_{l}(t)$, operating at a nominal pressure $p_{0}$ while subject to a time-varying heat load $Q_{h}(t)$. The TVS is composed of a vented branch and an injection loop. In the vented branch (blue line in Fig.7(a)), a mass flow rate $\dot{m}_{c}$ is extracted from the tank, expanded through a Joule-Thomson valve (1) and used as a cold source in a heat exchanger (2) before being expelled. In the injection loop, a mass flow rate $\dot{m}_{h}$ is circulated by a pump (3) into the hot side of the heat exchanger and reinjected into the tank as a subcooled liquid. This injection can be carried out via a submerged jet or a spray bar (Hastings et al., 2005; Wang et al., 2017; Hastings et al., 2003), but this distinction is unnecessary for the illustrative purposes of this work. An extensive literature review of TVS approaches is presented by Barsi (2011), who also introduced and tested two simplified approaches for the thermodynamic modelling of the problem (see Barsi and Kassemi 2013a and Barsi and Kassemi 2013b). Figure 7(b) maps the (ideal) key state of the fluids at the inlet and outlet of each component in the $T-s$ (temperature vs. specific entropy) diagram. The state of the liquid and the vapour in the tank are denoted with subscripts ($l$) and ($v$), the states ($1$) and ($2$) are the inlet and outlet of the heat exchanger on the vented line, while states ($1^{\prime}$) and ($2^{\prime}$) are their counterparts on the injection side.
@@
The control task is to keep the tank pressure below the maximum admissible one while venting the least amount of liquid. The simplest control problem, and also the one implemented in this work, consists of employing the pressure difference $\Delta p_{V}=p_{0}-p_{1}$ as the sole control parameter. The $\Delta p_{V}$ is linked to the mass flow rates in the J-T valve (1) and the pump (3) by limiting the operating range of the heat exchanger and by imposing that only latent heat is to be recovered from the venting line. Therefore, defining as $\Delta T_{l}(p_{1})=T_{l}-T_{1}(p_{1})$ the temperature drop produced by the J-T device and $\Delta h_{c}(p_{1})=h_{2}-h_{1}=h_{v,\text{sat}}(p_{1})-h_{l,\text{sat}}(p_{0})$ the specific latent heat remaining on the vented line, one has
@@
[…]
The simulation of the real environment and the digital twin are carried out using the same simplified model. In particular, we employ the homogeneous model presented in Barsi (2011), adapted to account for the venting line similarly to Mer et al. (2016a). This approach assumes that the liquid and the vapour are in saturation conditions. This is an oversimplification because usually the liquid is subcooled, and the vapour is superheated. We consider two control volumes: the saturated mixture and the solid wall. We refer to Marques et al. (2023) and Barsi and Kassemi (2013b) for more advanced models and to Panzarella and Kassemi (2003) and Panzarella et al. (2004) for an in-depth discussion on the limits of thermodynamic models. Focusing only on the storage problem (no out-flow except venting), the conservation of internal energy of the solid, and the mass and enthalpy of the saturated mixture read
@@
where $\rho$ is the density, $h$ is the specific enthalpy, and $\Delta h_{h}=h_{2^{\prime}}-h_{1^{\prime}}$ is the variation of specific enthalpy on the injection loop provided by the heat exchanger. All properties are at saturation conditions, and the subscripts $l$ and $v$ refer to the liquid and the vapour phases, respectively. We consider the mass $m_{s}$ and specific heat capacity $c_{s}$ of the wall to be constant. Denoting as $H=\rho h$ the (volumetric) enthalpies and using the chain rule to have $d(\cdot)/dt=(d(\cdot)/dp)\;(dp/dt)$, these can be written as
@@
The heat transferred from the wall to the fluid mixture is modelled using Newton’s cooling law $Q_{s\rightarrow f}=U_{s}(T_{s}-T)$, where $U_{s}$ is the overall heat transfer coefficient. Taking all liquid and vapour properties from thermophysical property libraries, this is a closed set of implicit differential equations for $\bm{s}=[T_{s},V_{l},p]$. The fluid properties are retrieved through the open-source library CoolProp (Lemmon et al., 2018), which is based on multiparameter Helmholtz-energy-explicit-type formulations and also provides accurate derivatives.
@@
Model predictions are thus possible once these three parameters are provided. We consider these the model closure for the digital twin, hence $\bm{w}_{p}=[U_{s},U_{HX},K_{JT}]$. These are to be inferred from real-time data by minimizing the following cost function.

### v4

@@ 5.1 Wind Turbine Control
The wind energy community has put significant effort into the development of digital twins of onshore and offshore wind turbines, with applications ranging from load estimations [Pimenta et al., 2020, Branlard et al., 2023] to real-time condition monitoring [Olatunji et al., 2021, Fahim et al., 2022]. In this work, we solely focus on the rotor’s aerodynamic performance and the twinning process seeks to identify a closure law for the power coefficient from data, starting from (25) and (26). In fact, the BEMT, commonly used to obtain the $C_{p}$ curve, assumes that the flow is steady, and this is questionable in the presence of large fluctuations in the wind speed [Leishman, 2002]. Moreover, the coefficient can change over time as a result of turbine wear (e.g. due to blade erosion) or in the presence of extreme conditions (e.g. ice, dirt or bug buildup) [Staffell and Green, 2014]. This is why any strategy based on a pre-computed $C_{p}$ curve is likely sub-optimal [Johnson, 2004, 2006] and the idea of continuous updates of this closure law from operational data has merit. Although this exercise is purely illustrative, the approach can be readily extended to more complex and multi-disciplinary model formulations.
The digital twin employs the power coefficient parametrization formulated by Saint-Drenan et al. [2020]
@@
$\mathcal{J}_{p}=\frac{1}{2}\int_{0}^{T}\Big{(}\omega_{g}(t;\bm{w}_{p})-\check{\omega}_{g}(t)\Big{)}^{2}\;\mathrm{dt}\,.$ (28)
@@ 5.2 Trajectory control of flapping wings micro air vehicles
This test case considers the control of a Flapping-Wing Micro Air Vehicle (FWMAV) which aims to move from one position to another in the shortest possible time. FWMAVs are sometimes considered better alternatives to fixed-wing and rotary-wing micro aerial vehicles because of their higher agility and manoeuvrability [Haider et al., 2020]. The design and control of these drones seek to mimic the remarkable flight performances of hummingbirds, which can undertake complex manoeuvres with an impressive response time (see, for example, Cheng et al. 2016 and Ortega-Jiménez and Dudley 2018).
Reproducing these performances requires advanced knowledge of the flapping wing’s aerodynamics in the low Reynolds number regime, as well as the fluid-structure interaction resulting from the wing’s flexibility. These remain significant challenges for advanced numerical simulations [Fei et al., 2019, Xue et al., 2023]. The proposed test case shows how simplified models of the wing’s aerodynamics could be derived in real-time while an FWAVs seeks to achieve its control target. The considered problem is illustrated in Figure 6(a). A drone initially at $\bm{x}_{0}=[x_{0},z_{0}]$ m and $\bm{\dot{x}}_{0}=0$ m/s is requested to move to $\bm{\tilde{x}_{f}}=[\tilde{x}_{f},\tilde{z}_{f}]$ m as fast as possible, and then hover in that position (i.e. $\dot{\bm{\tilde{x}}}_{f}=[0,0]$) by continuously adapting its flapping wing motion. We challenge the controller by introducing gusty wind disturbances $\bm{u}_{\infty}$,
@@
The body frame is first pitched by a stroke plane angle $\beta_{w}(t)$ along $y^{\prime}$ to define the stroke plane frame ($x^{\prime\prime},y^{\prime\prime},z^{\prime\prime}$). The wing tips are constrained to flap in the stroke plane ($x^{\prime\prime},y^{\prime\prime}$) as it flaps along the normal $z^{\prime\prime}$. The flapping frame ($x^{\prime\prime\prime},y^{\prime\prime\prime},z^{\prime\prime\prime}$) follows this rotational motion and defines the flapping angle $\phi(t)$ between the wing symmetry axis $y^{\prime\prime\prime}$ and the lateral direction $y^{\prime\prime}$ (Figure 6(b)). The last frame is the wing-fixed frame $(\xi,\gamma,\zeta)$ that results from the pitching rotation along the wing symmetry axis. The pitching angle $\alpha(t)$ is then defined between the chord-normal direction $\xi$ and the stroke plane axis $x^{\prime\prime\prime}$. The reader is referred to Whitney and Wood [2010] and Cai et al. [2021] for a complete treatment of the flapping wing kinematics. We select an harmonic parametrization for $\phi$ and $\alpha$:
@@
We consider a drone with mass $m_{b}=3$ g and a constant flapping frequency of $f=20$ Hz. Given the initial position $\bm{x}_{o}=[0,0]$ m, we set the target position at $\bm{x}_{f}=[5,5]$ m. We assume that the manoeuvre should take less than $n_{t}=100$ flapping cycles, hence we consider an observation time of $T_{o}=5$ s. This defines the observation time for the twinning. We fix the pitching amplitude to $A_{\alpha}=45$ deg and let the flapping amplitude vary in the range $A_{\phi}\in[50;88]$ deg. These values are comparable to what is observed in hummingbird’s flight, although the harmonic parameterization in (29) oversimplifies their actual wing dynamics [Kruyt et al., 2014]. The stroke plane angle is bounded in the range $\beta_{w}\in[-30,30]$ deg.
We assume that the drone’s wings are semi-elliptical and rigid with a span $R=0.05$ m and a mean chord $\overline{c}=0.01$ m. The wing roots are offset of $R_{0}=0.0225$ m from the body barycenter, which is also the centre of rotation of the wings. We decouple their motion from the body such that the wing inertia does not influence the body motion [Taha et al., 2012]. This hypothesis simplifies the body dynamics, which becomes only a function of the aerodynamic forces produced by the wings and the gravitational force. These are assumed to apply on the barycenter of the body, which is treated as a material point. Therefore, restricting the dynamics to the $(x,z)$ plane, the force balance gives
@@
where $\rho$ is the air density, $S_{b}=0.0005$ m2 is the body’s cross-sectional area, $C_{D,b}$ is the body’s drag coefficient, here taken as unitary, and $U_{b}$ is the modulus of the relative velocity between the drone body and the wind, i.e. $U_{b}=\sqrt{(\dot{x}+u_{\infty})^{2}+\dot{z}^{2}}$. The wind disturbance $u_{\infty}$ results from a stationary Gaussian process with mean value $\overline{u}_{\infty}=1$ m/s, in opposite direction to the $x$ versor, and covariance matrix defined by a Gaussian kernel $\kappa(u_{\infty}(t),u_{\infty}(t+\tau))=\exp(-\tau^{2}/(2\sigma^{2}_{f}))$ with $\sigma_{f}=3\mbox{s}$.
To compute the wing forces in (34) and (35) we follow the semi-empirical quasi-steady approach of Lee et al. [2016], based on the work by Dickinson et al. [1999]. This is a common approach in control-oriented investigations (see e.g Cai et al. [2021], Fei et al. [2019]) because it provides a reasonable compromise between accuracy and computational cost, at least for the case of smooth flapping kinematics and near hovering conditions (see Lee et al. [2016]).
The approach is based on BEMT, which partitions the wing into infinitesimal elements (see Figure 7(a)). Each element produces an infinitesimal aerodynamic force that must be integrated along the span to get the total lift and drag. For smooth flapping motion (29), the forces due to the flapping dominate all the other physical phenomena (Lee et al. [2016]) so that the lift $L_{w}$ and drag $D_{w}$ expressions are
@@
The force coefficients $C_{L,w}$ and $C_{D,w}$ in equations (37) and (38) are modeled similarly to Lee et al. [2016] and Sane and Dickinson [2001]:
@@
where $\bm{w}_{p}=[a,b,c]$ are the closure parameters for this problem. This parametrization approximates the influence of the Leading Edge Vortex (LEV), generated at a high pitching angle due to flow separation and stably attached on the suction side (see Sane [2003]). The corresponding force coefficients depend on the effective angle of attack $\alpha_{e}=\cos^{-1}(U_{w,z}$/$||\bm{U_{w}}||_{2})$, defined between the chord direction $\zeta$ and the relative velocity of the wing $\bm{U_{w}}$.
@@
In this test case, the real environment is simulated using $\bm{w_{p}}=[a,b,c]=[1.71,0.043,1.595]$ in (37) and (38), as approximated from [Lee et al., 2016] for semi-elliptical wings. For the digital twin, these are the parameters to be identified while interacting with the environment to minimize the cost function:
@@ 5.3 Thermal Management of Cryogenic Storage
We consider the thermal management of a cryogenic tank. This is a fundamental challenge in long-duration space missions, which require storing large amounts of cryogenic propellant for long periods. These fluids must be stored in the liquid phase to maximize the volumetric energy density. However, their storage temperatures are extremely low (e.g., -250 C for liquid hydrogen), and their latent heat of evaporation is much smaller than that of non-cryogenic propellants. These combined factors challenge the storage because evaporation due to thermal loads produces a continuous pressure rise over time [Salzman, 1996, Motil et al., 2007, Chai and Wilhite, 2014].
@@
The efficient management of this problem requires a combination of advanced insulation strategies (see Mer et al. 2016b, Jiang et al. 2021) and venting control strategies [Lin et al., 1991]. In this test case, we consider a classic active Thermodynamic Venting System (TVS), similar to the one presented in Lin et al. [1991] and recently characterized in Imai et al. [2020], Qin et al. [2021]. Figure 8(a) shows the schematics of the TVS configuration.
@@
Figure 8: Schematic of the thermal management problem considered in this work. A cryogenic tank of volume $V_{t}$, containing a volume $V_{l}$ of saturated liquid and $V_{t}-V_{l}$ saturated vapour, is subject to a heat load $Q_{h}(t)$. The figure on the right traces the state of the system on the temperature-specific entropy diagram considering the simplified model used to simulate the environment.
@@
We consider a tank of volume $V_{t}$ filled with a volume of liquid $V_{l}(t)$, operating at a nominal pressure $p_{0}$ while subject to a time-varying heat load $Q_{h}(t)$. The TVS is composed of a vented branch and an injection loop. In the vented branch (blue line in Fig.8(a)), a mass flow rate $\dot{m}_{c}$ is extracted from the tank, expanded through a Joule-Thomson valve (1) and used as a cold source in a heat exchanger (2) before being expelled. In the injection loop, a mass flow rate $\dot{m}_{h}$ is circulated by a pump (3) into the hot side of the heat exchanger and reinjected into the tank as a subcooled liquid. This injection can be carried out via a submerged jet or a spray bar [Hastings et al., 2005, Wang et al., 2017, Hastings et al., 2003], but this distinction is unnecessary for the illustrative purposes of this work. An extensive literature review of TVS approaches is presented by Barsi [2011], who also introduced and tested two simplified approaches for the thermodynamic modelling of the problem (see Barsi and Kassemi 2013a and Barsi and Kassemi 2013b). Figure 8(b) maps the (ideal) key state of the fluids at the inlet and outlet of each component in the $T-s$ (temperature vs. specific entropy) diagram. The state of the liquid and the vapour in the tank are denoted with subscripts ($l$) and ($v$), the states ($1$) and ($2$) are the inlet and outlet of the heat exchanger on the vented line, while states ($1^{\prime}$) and ($2^{\prime}$) are their counterparts on the injection side.
@@
The control task is to keep the tank pressure below the maximum admissible one while venting the least amount of liquid. The simplest control problem, implemented in this work, consists of employing the pressure difference $\Delta p_{V}=p_{0}-p_{1}$ as the sole control parameter. The $\Delta p_{V}$ is linked to the mass flow rates in the J-T valve (1) and the pump (3) by limiting the operating range of the heat exchanger and by imposing that only latent heat is to be recovered from the venting line. Therefore, defining as $\Delta T_{l}(p_{1})=T_{l}-T_{1}(p_{1})$ the temperature drop produced by the J-T device and $\Delta h_{c}(p_{1})=h_{2}-h_{1}=h_{v,\text{sat}}(p_{1})-h_{l,\text{sat}}(p_{0})$ the specific latent heat remaining on the vented line, one has
@@
We consider a cryogenic tank with volume $V_{t}=60$ m3 with a desired operational pressure $\tilde{p}=500$ kPa filled with liquid hydrogen. At the start of each episode, the tank is $70\%$ full and pressurized to an initial value sampled from a normal distribution with mean $\overline{p}_{0}=500$ kPa and standard deviation $\sigma_{p_{0}}=25$ kPa. The thermal load applied to the tank fluctuates from a minimum of roughly $50$ W to a maximum of $300$ W with a period of $T_{h}\approx 1.5$ h. This cyclic load mimics the large thermal fluctuations a tank could face in orbit as it is periodically exposed to direct sun radiation. The chosen period is approximately the time it takes the International Space Station (ISS) to complete an orbit around Earth. The load profile is shown in Figure 9(a). This load is constructed as a pseudo-random signal made of Gaussian-like peaks randomly varying within $[0.9,1.1]$ times the nominal peak, lifted by the minimal load and polluted by uniform noise with a standard deviation equal to $10\%$ the nominal peak.
The simulation of the real environment and the digital twin are carried out using the same simplified model. In particular, we employ the homogeneous model presented in Barsi [2011], adapted to account for the venting line similarly to Mer et al. [2016a]. This approach assumes that the liquid and the vapour are in saturation conditions. This is an oversimplification because the liquid is typically subcooled and the vapor is usually superheated. We refer to Marques et al. [2023] and Barsi and Kassemi [2013b] for more advanced models and to Panzarella and Kassemi [2003] and Panzarella et al. [2004] for an in-depth discussion on the limits of thermodynamic models. Focusing only on the storage problem (no out-flow except venting), the conservation of internal energy of the solid, and the mass and enthalpy of the saturated mixture read
@@
where $\rho$ is the density, $h$ is the specific enthalpy, and $\Delta h_{h}=h_{2^{\prime}}-h_{1^{\prime}}$ is the variation of specific enthalpy on the injection loop provided by the heat exchanger. All properties are at saturation conditions, and the subscripts $l$ and $v$ refer to the liquid and the vapour phases, respectively. We consider the mass $m_{s}$ and the wall’s specific heat capacity $c_{s}$ constant. Denoting as $H=\rho h$ the (volumetric) enthalpies and using the chain rule to have $d(\cdot)/dt=(d(\cdot)/dp)\;(dp/dt)$, these can be written as
@@
The heat transferred from the wall to the fluid mixture is modelled using Newton’s cooling law $Q_{s\rightarrow f}=U_{s}(T_{s}-T)$, where $U_{s}$ is the overall heat transfer coefficient. Taking all liquid and vapour properties from thermophysical property libraries, this is a closed set of implicit differential equations for $\bm{s}=[T_{s},V_{l},p]$. The fluid properties are retrieved through the open-source library CoolProp [Lemmon et al., 2018], which is based on multiparameter Helmholtz-energy-explicit-type formulations and also provides accurate derivatives.
@@
Model predictions are thus possible once these three parameters are provided. We consider these the model closure for the digital twin, hence $\bm{p}=\bm{w}_{p}=[U_{s},U_{HX},K_{JT}]$. These are to be inferred from real-time data by minimizing the following cost function.